OpenStudy (anonymous):

Anyone??? What term do you need to add to each side of 6x^2 + 24x = 1 to complete the square? Choices: 4 1/4 12 1/12

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

6x^2 + 24x = 1 factor out the left 6(x^2+4x) = 1 complete the square inside now by adding the square of half the x part 6(x^2+4x+4-4) = 1 6(x+2)^2 -6(4) = 1

OpenStudy (amistre64):

its either 4 or 12 ... the choices seem odd to me

OpenStudy (amistre64):

in effect, your adding 24 to each side :/

OpenStudy (anonymous):

yea i see what your saying!! math.. gotta love it

OpenStudy (amistre64):

the answer is most likely 4 that they want; but they word the question poorly

OpenStudy (anonymous):

the website that im studying on says its 4 = /

OpenStudy (amistre64):

what happens when we add 4 to each side? 6x^2 + 24x +4= 1+4 6x^2 + 24x +4 = 5 6x^2 +24x -1 = 0 if this is a perfect square; we should end up with one root for x x = -6+- sqrt(36-4(-6))/12 x = -6+- sqrt(60)/12 aint no way; we still got a value under the sqrt

OpenStudy (anonymous):

ugh ok another one to have answered tomorrow... thanks for trying though im writing all this down

OpenStudy (anonymous):

4

OpenStudy (anonymous):

(x + 4x) + 4 = (x + 2)^2

OpenStudy (anonymous):

It seems the question just focus on what number adding to complete square, not the result!

OpenStudy (anonymous):

oh i see thank you so much!

OpenStudy (anonymous):

I missed square, (x^2 + 4x) + 4 = (x + 2)^2

OpenStudy (anonymous):

Thanks!!

OpenStudy (amistre64):

yes, they are focusing on that while asking for " What term do you need to add to each side" which is NOT that ...